# Reinforced Generation of Combinatorial Structures: Hardness of Approximation

### v5

@@ Appendix D Proof of the Hardness of Approximation for TSP
We will construct collections of edges $T^{=},T^{\neq}$, and $T_{1}^{=}$ from $E^{=},E^{\neq}$, $E_{1}^{=}$, and set $T=T^{=}\cup T^{\neq}\cup T_{1}^{=}$. First let us specify how we construct $T^{=}$ and $T^{\neq}$.
  • We add a single copy of each (forced) edge in $T^{\neq}$. Note that there are two parallel forced edges corresponding to every inequality constraint in $I^{\neq}$; we add a single copy of both.
  • For a variable set to $1$, $T^{=}$ contains exactly one copy of both unforced edges in $E^{=}$ adjacent to the corresponding vertex in $G$.
Note that because every $I^{=}$ and $I^{\neq}$ clause is satisfied, for each $i\in[n]$ there is a single $b^{*}\in\{0,1\}$ such that $x_{i,j,b}=\mathbf{1}\{b=b^{*}\}$. So $T^{=}$ contains exactly half the edges in $E^{=}$. We can hence calculate $w(T^{\neq})=20nk$, and $w(T^{=})=11nk$. Next we describe how to construct $T_{1}^{=}$.
For each distinct clause of the form $P^{=}_{1}(x_{i_{1},j_{1},b_{1}},x_{i_{2},j_{2},b_{2}},x_{i_{3},j_{3},b_{3}})$, we consider the optimal spanning tour $Q^{*}=\arg\min_{Q\in\mathbb{Q}_{\mathbf{z}}}w(T)$, where $\mathbf{z}=(x_{i_{1},j_{1},b_{1}},x_{i_{2},j_{2},b_{2}},x_{i_{3},j_{3},b_{3}})$. We will add all non-special edges of $Q^{*}$ on the corresponding contact vertices $\{x_{i_{1},j_{1},b_{1}},x_{i_{2},j_{2},b_{2}},x_{i_{3},j_{3},b_{3}}\}$, central vertex $s$, and auxiliary vertices $a_{1},\ldots,a_{n_{\textup{aux}}}$ specific to the clause.
@@
$c(H)-(x_{i_{1},j_{1},b_{1}}+x_{i_{2},j_{2},b_{2}}+x_{i_{3},j_{3},b_{3}})+2w(H)\cdot\left(\mathbf{1}\{11k\in\{j_{1},j_{2},j_{3}\}\}+\mathbf{1}\{P^{=}_{1}(x_{i_{1},j_{1},b_{1}},x_{i_{2},j_{2},b_{2}},x_{i_{3},j_{3},b_{3}})=0\}\right),$
@@
Summing over all clauses in $I_{1}^{=}$, we get
$w(T_{1}^{=})\leq c(H)\cdot 2nk/3-nk+2w(H)\cdot(2n+\Delta),$
where we used that the number of $I_{1}^{=}$ clauses is $2nk/3$, the fact that at most $2n$ clauses satisfy $11k\in\{j_{1},j_{2},j_{3}\}$, and at most $\Delta$ clauses are unsatisfied.
@@
The number of edges of $T^{=}$ and $T^{\neq}$ adjacent to each vertex in $G$ is even by construction.
Every non-contact vertex is also adjacent to an even number of edges in $T_{1}^{=}$. Finally, since in every spanning tour in $\mathbb{Q}_{\mathbf{z}}$ for $\mathbf{z}\in\{0,1\}^{3}$, each special edge (which we abstained from adding to $T_{1}^{=}$) is used an even number of times, it also holds that every contact vertex is adjacent to an even number of edges in $T_{1}^{=}$. ∎
@@
  • The contact vertices $x_{i,j,b}$ are connected to $s$ using only edges in $T_{1}^{=}$ if either (1) $x_{i,j,b}=0$, or (2) $j=11k$.
  • All checker vertices, and contact vertices set to $1$ are connected to either $x_{i,11k,0}$ or $x_{i,11k,1}$ by edges in $T^{=}$ and $T^{\neq}$. Therefore they are also connected to $s$ in $T$ by the previous step.
@@
Let us write $T=T^{=}\cup T^{\neq}\cup T_{1}^{=}$, where $T^{=},T^{\neq}$, and $T_{1}^{=}$ are the edges in $T$ belonging to $^{=}, EE^{\neq}$, and $E_{1}^{=}$ respectively. We define $U^{=}$ to be $T^{=}$ with all edges appearing twice removed.
We will assign $x_{i,j,b}$ based on the “local” view of the variable $x_{i,j,b}$ as a vertex in $U^{=}$. In particular, let $l_{i,j,b}$ and $r_{i,j,b}$ denote the number of copies of the edges $(x_{i,j-1,b},x_{i,j,b})$ and $(x_{i,j,b},x_{i,j+1,b})$ in $T^{=}$ respectively. We say that the variable $x_{i,j,b}$ is “honest” if $l_{i,j,b}=r_{i,j,b}$, and “dishonest” otherwise. We will always assign an honest variable $x_{i,j,b}=l_{i,j,b}$. Note that this is equivalent to the assignment $x_{i,j,b}=r_{i,j,b}$.
@@
  • For each inequality clause $P^{\neq}(x_{i,j,0},x_{i,M(j),1})$ in $I^{\neq}$, note that the variables involved are either both honest, or both dishonest. In the case that both are dishonest, we assign each variable a random value in $\{0,1\}$, conditioned on $x_{i,j,0}\neq x_{i,M(j),1}$.
  • For each $3\textup{LIN}({2})$ clause $P^{=}_{1}(x_{i_{1},j_{1},b_{1}},x_{i_{2},j_{2},b_{2}},x_{i_{3},j_{3},b_{3}})$ in $I_{1}^{=}$ containing at least one dishonest variable, we arbitrarily set the dishonest variables involved, while ensuring that the clause is satisfied. Note that this is always possible as long as there is at least one such dishonest variable.
@@
  • $p_{ch}$ denotes the number of inequality constraints in $I^{\neq}$ such that both involved checkers are honest and assigned the same value.
  • $r$ is the number of unsatisfied clauses in $I_{1}^{=}$ where all involved variables are honest.
@@
  • Inequality clauses: An inequality clause in $I^{\neq}$ is unsatisfied when both checkers it involves are honest and assigned the same value. So, the number of violated inequality clauses is equal to $p_{ch}$.
  • 3LIN(2) clauses: By definition, all clauses in $I_{1}^{=}$ involving dishonest variables are satisfied. So the number of unsatisfied clauses in $I_{1}^{=}$ is equal to $r$.
Since the instance $I$ is defined as $I^{=}+2I^{\neq}+2I_{1}^{=}$, the above bounds imply that the number of unsatisfied clauses in $I$ is at most $d_{ch}+d_{co}+2p_{ch}+2r$. ∎
@@
We will separately bound the weight of edges in $^{=}, TT^{\neq}$, and $T_{1}^{=}$. Let us denote by $t_{co}$ the number of honest contact variables assigned $1$. For $b\in\{0,1\}$ let $p_{ch}^{b}$ denote the number of inequality clauses such that both involved checkers are assigned $b$. Note that $p_{ch}^{0}+p_{ch}^{1}=p_{ch}$. Finally, let $u_{co}$ be the number of connected components in $U^{=}\cup T^{\neq}\cup T_{1}^{=}$ containing only edges from $E_{1}^{=}$.
  • Bounding $U^{=}$: By double-counting, we can write
    $w(U^{=})=\frac{1}{2}\sum_{i,j,b}\left(l_{i,j,b}+r_{i,j,b}\right)=\frac{1}{2}\left(\sum_{i,j:j\mid 11,b}\left(l_{i,j,b}+r_{i,j,b}\right)+\sum_{i,j:j\nmid 11}v_{i,j}\right),$
@@
    Next, we will bound the second sum. We have that the sum of degrees of $x_{i,j,0}$ and $x_{i,M(j),1}$ in $U^{=}$ is even; equivalently, $v_{i,j}$ must be even. So, $v_{i,j}\geq 2$ unless both involved checkers are assigned 0. Furthermore, we have $v_{i,j}\geq 4$ if both involved checkers are honest and assigned $1$.
@@
    Together, these bounds imply $w(U^{=})\geq 10nk+d_{co}/2+t_{co}+p_{ch}^{1}-p_{ch}^{0}$.
  • Bounding $T^{=}\setminus U^{=}$: Since $T=T^{=}\cup T^{\neq}\cup T_{1}^{=}$ is a spanning tour of $G$, we have that the total weight of edges in $T^{=}\setminus U^{=}$ is at least twice the number of connected components of $U^{=}\cup T^{\neq}\cup T_{1}^{=}$. So, $w(T^{=}\setminus U^{=})\geq 2p_{ch}^{0}+2u_{co}$.
@@
    $w(T^{\neq})=\sum_{i\in[n],j\in[11k]:j\nmid 11}\kappa_{i,j}.$
    Since $T$ was a valid tour, we have $\kappa_{i,j}\geq 2$ for all such $i,j$. Furthermore, if either of the involved vertices is dishonest, the fact that the degree of that vertex in $T$ is even implies that $\kappa_{i,j}$ is odd. As a consequence, we have $\kappa_{i,j}\geq 3$ for at least $d_{ch}/2$ such $(i,j)$. This implies $w(T^{\neq})\geq 2\cdot 10nk+d_{ch}/2$.
  • 3LIN(2) clauses: Let $\zeta=P^{=}_{1}(x_{i_{1},j_{1},b_{1}},x_{i_{2},j_{2},b_{2}},x_{i_{3},j_{3},b_{3}})$ be a clause in $I_{1}^{=}$. Let $T_{\zeta}$ be the set of edges in $T_{1}^{=}$ involved in clause $\zeta$. We will extract a valid spanning tour $Q$ of $H$ from $T_{\zeta}$, and then use the soundness property of $H$ to argue that $w(T_{\zeta})$ is large.
@@
    Summing over all $2nk/3$ clauses $c$ in $I_{1}^{=}$, we get
    $w(T_{1}^{=})\geq(2nk/3)\cdot s(H)-2u_{co}-t_{co}+r,$
@@
Together, these imply the lower bound $w(T)=w(U^{=})+w(T^{=})+w(T^{\neq})+w(T_{1}^{=})\geq nk\cdot(30+2\cdot s(H)/3)+(d_{ch}+d_{co})/2+p_{ch}+r$. ∎
@@ D.2 Description of our gadget
We note that this is in alignment with the discussion in Section 4; every satisfying assignment to $P^{=}_{1}$ admits a tour of cost at most $c(H)=10$, whereas for every tour $Q$ corresponding to an unsatisfying assignment, we have $k_{1}^{H}(Q)=0$ and $k_{2}^{H}(Q)$ is even; hence $Q$ has weight at least $s(H)+\mathbf{1}\{k_{1}^{H}(Q)=0\}\cdot\mathbf{1}\{k_{2}^{H}(Q)\text{ is even}\}=11$. We can now prove Theorem 4.1.

### v6

@@ Appendix D Proof of the Hardness of Approximation for TSP
We will construct collections of edges $\texttt{T}^{=}_{2},\texttt{T}^{\neq}_{2}$, and $\texttt{T}^{\equiv 1}_{3,2}$ from $\texttt{E}^{=}_{2},\texttt{E}^{\neq}_{2}$, $\texttt{E}^{\equiv 1}_{3,2}$, and set $T=\texttt{T}^{=}_{2}\cup\texttt{T}^{\neq}_{2}\cup\texttt{T}^{\equiv 1}_{3,2}$. First let us specify how we construct $\texttt{T}^{=}_{2}$ and $\texttt{T}^{\neq}_{2}$.
  • We add a single copy of each (forced) edge in $\texttt{T}^{\neq}_{2}$. Note that there are two parallel forced edges corresponding to every inequality constraint in $\texttt{I}^{\neq}_{2}$; we add a single copy of both.
  • For a variable set to $1$, $\texttt{T}^{=}_{2}$ contains exactly one copy of both unforced edges in $\texttt{E}^{=}_{2}$ adjacent to the corresponding vertex in $G$.
Note that because every $\texttt{I}^{=}_{2}$ and $\texttt{I}^{\neq}_{2}$ clause is satisfied, for each $i\in[n]$ there is a single $b^{*}\in\{0,1\}$ such that $x_{i,j,b}=\mathbf{1}\{b=b^{*}\}$. So $\texttt{T}^{=}_{2}$ contains exactly half the edges in $\texttt{E}^{=}_{2}$. We can hence calculate $w(\texttt{T}^{\neq}_{2})=20nk$, and $w(\texttt{T}^{=}_{2})=11nk$. Next we describe how to construct $\texttt{T}^{\equiv 1}_{3,2}$.
For each distinct clause of the form $\texttt{P}^{\equiv 1}_{3,2}(x_{i_{1},j_{1},b_{1}},x_{i_{2},j_{2},b_{2}},x_{i_{3},j_{3},b_{3}})$, we consider the optimal spanning tour $Q^{*}=\arg\min_{Q\in\mathbb{Q}_{\mathbf{z}}}w(T)$, where $\mathbf{z}=(x_{i_{1},j_{1},b_{1}},x_{i_{2},j_{2},b_{2}},x_{i_{3},j_{3},b_{3}})$. We will add all non-special edges of $Q^{*}$ on the corresponding contact vertices $\{x_{i_{1},j_{1},b_{1}},x_{i_{2},j_{2},b_{2}},x_{i_{3},j_{3},b_{3}}\}$, central vertex $s$, and auxiliary vertices $a_{1},\ldots,a_{n_{\textup{aux}}}$ specific to the clause.
@@
$c(H)-(x_{i_{1},j_{1},b_{1}}+x_{i_{2},j_{2},b_{2}}+x_{i_{3},j_{3},b_{3}})+2w(H)\cdot\left(\mathbf{1}\{11k\in\{j_{1},j_{2},j_{3}\}\}+\mathbf{1}\{\texttt{P}^{\equiv 1}_{3,2}(x_{i_{1},j_{1},b_{1}},x_{i_{2},j_{2},b_{2}},x_{i_{3},j_{3},b_{3}})=0\}\right),$
@@
Summing over all clauses in $\texttt{I}^{\equiv 1}_{3,2}$, we get
$w(\texttt{T}^{\equiv 1}_{3,2})\leq c(H)\cdot 2nk/3-nk+2w(H)\cdot(2n+\Delta),$
where we used that the number of $\texttt{I}^{\equiv 1}_{3,2}$ clauses is $2nk/3$, the fact that at most $2n$ clauses satisfy $11k\in\{j_{1},j_{2},j_{3}\}$, and at most $\Delta$ clauses are unsatisfied.
@@
The number of edges of $\texttt{T}^{=}_{2}$ and $\texttt{T}^{\neq}_{2}$ adjacent to each vertex in $G$ is even by construction.
Every non-contact vertex is also adjacent to an even number of edges in $\texttt{T}^{\equiv 1}_{3,2}$. Finally, since in every spanning tour in $\mathbb{Q}_{\mathbf{z}}$ for $\mathbf{z}\in\{0,1\}^{3}$, each special edge (which we abstained from adding to $\texttt{T}^{\equiv 1}_{3,2}$) is used an even number of times, it also holds that every contact vertex is adjacent to an even number of edges in $\texttt{T}^{\equiv 1}_{3,2}$. ∎
@@
  • The contact vertices $x_{i,j,b}$ are connected to $s$ using only edges in $\texttt{T}^{\equiv 1}_{3,2}$ if either (1) $x_{i,j,b}=0$, or (2) $j=11k$.
  • All checker vertices, and contact vertices set to $1$ are connected to either $x_{i,11k,0}$ or $x_{i,11k,1}$ by edges in $\texttt{T}^{=}_{2}$ and $\texttt{T}^{\neq}_{2}$. Therefore they are also connected to $s$ in $T$ by the previous step.
@@
Let us write $T=\texttt{T}^{=}_{2}\cup\texttt{T}^{\neq}_{2}\cup\texttt{T}^{\equiv 1}_{3,2}$, where $\texttt{T}^{=}_{2},\texttt{T}^{\neq}_{2}$, and $\texttt{T}^{\equiv 1}_{3,2}$ are the edges in $T$ belonging to $\texttt{E}^{=}_{2}$, $\texttt{E}^{\neq}_{2}$, and $\texttt{E}^{\equiv 1}_{3,2}$ respectively. We define $\texttt{U}^{=}_{2}$ to be $\texttt{T}^{=}_{2}$ with all edges appearing twice removed.
We will assign $x_{i,j,b}$ based on the “local” view of the variable $x_{i,j,b}$ as a vertex in $\texttt{U}^{=}_{2}$. In particular, let $l_{i,j,b}$ and $r_{i,j,b}$ denote the number of copies of the edges $(x_{i,j-1,b},x_{i,j,b})$ and $(x_{i,j,b},x_{i,j+1,b})$ in $\texttt{T}^{=}_{2}$ respectively. We say that the variable $x_{i,j,b}$ is “honest” if $l_{i,j,b}=r_{i,j,b}$, and “dishonest” otherwise. We will always assign an honest variable $x_{i,j,b}=l_{i,j,b}$. Note that this is equivalent to the assignment $x_{i,j,b}=r_{i,j,b}$.
@@
  • For each inequality clause $\texttt{P}^{\neq}_{2}(x_{i,j,0},x_{i,M(j),1})$ in $\texttt{I}^{\neq}_{2}$, note that the variables involved are either both honest, or both dishonest. In the case that both are dishonest, we assign each variable a random value in $\{0,1\}$, conditioned on $x_{i,j,0}\neq x_{i,M(j),1}$.
  • For each $3\textup{LIN}({2})$ clause $\texttt{P}^{\equiv 1}_{3,2}(x_{i_{1},j_{1},b_{1}},x_{i_{2},j_{2},b_{2}},x_{i_{3},j_{3},b_{3}})$ in $\texttt{I}^{\equiv 1}_{3,2}$ containing at least one dishonest variable, we arbitrarily set the dishonest variables involved, while ensuring that the clause is satisfied. Note that this is always possible as long as there is at least one such dishonest variable.
@@
  • $p_{ch}$ denotes the number of inequality constraints in $\texttt{I}^{\neq}_{2}$ such that both involved checkers are honest and assigned the same value.
  • $r$ is the number of unsatisfied clauses in $\texttt{I}^{\equiv 1}_{3,2}$ where all involved variables are honest.
@@
  • Inequality clauses: An inequality clause in $\texttt{I}^{\neq}_{2}$ is unsatisfied when both checkers it involves are honest and assigned the same value. So, the number of violated inequality clauses is equal to $p_{ch}$.
  • 3LIN(2) clauses: By definition, all clauses in $\texttt{I}^{\equiv 1}_{3,2}$ involving dishonest variables are satisfied. So the number of unsatisfied clauses in $\texttt{I}^{\equiv 1}_{3,2}$ is equal to $r$.
Since the instance I is defined as $\texttt{I}^{=}_{2}+2\texttt{I}^{\neq}_{2}+2\texttt{I}^{\equiv 1}_{3,2}$, the above bounds imply that the number of unsatisfied clauses in I is at most $d_{ch}+d_{co}+2p_{ch}+2r$. ∎
@@
We will separately bound the weight of edges in $\texttt{T}^{=}_{2}$, $\texttt{T}^{\neq}_{2}$, and $\texttt{T}^{\equiv 1}_{3,2}$. Let us denote by $t_{co}$ the number of honest contact variables assigned $1$. For $b\in\{0,1\}$ let $p_{ch}^{b}$ denote the number of inequality clauses such that both involved checkers are assigned $b$. Note that $p_{ch}^{0}+p_{ch}^{1}=p_{ch}$. Finally, let $u_{co}$ be the number of connected components in $\texttt{U}^{=}_{2}\cup\texttt{T}^{\neq}_{2}\cup\texttt{T}^{\equiv 1}_{3,2}$ containing only edges from $\texttt{E}^{\equiv 1}_{3,2}$.
  • Bounding $\texttt{U}^{=}_{2}$: By double-counting, we can write
    $w(\texttt{U}^{=}_{2})=\frac{1}{2}\sum_{i,j,b}\left(l_{i,j,b}+r_{i,j,b}\right)=\frac{1}{2}\left(\sum_{i,j:j\mid 11,b}\left(l_{i,j,b}+r_{i,j,b}\right)+\sum_{i,j:j\nmid 11}v_{i,j}\right),$
@@
    Next, we will bound the second sum. We have that the sum of degrees of $x_{i,j,0}$ and $x_{i,M(j),1}$ in $\texttt{U}^{=}_{2}$ is even; equivalently, $v_{i,j}$ must be even. So, $v_{i,j}\geq 2$ unless both involved checkers are assigned 0. Furthermore, we have $v_{i,j}\geq 4$ if both involved checkers are honest and assigned $1$.
@@
    Together, these bounds imply $w(\texttt{U}^{=}_{2})\geq 10nk+d_{co}/2+t_{co}+p_{ch}^{1}-p_{ch}^{0}$.
  • Bounding $\texttt{T}^{=}_{2}\setminus\texttt{U}^{=}_{2}$: Since $T=\texttt{T}^{=}_{2}\cup\texttt{T}^{\neq}_{2}\cup\texttt{T}^{\equiv 1}_{3,2}$ is a spanning tour of $G$, we have that the total weight of edges in $\texttt{T}^{=}_{2}\setminus\texttt{U}^{=}_{2}$ is at least twice the number of connected components of $\texttt{U}^{=}_{2}\cup\texttt{T}^{\neq}_{2}\cup\texttt{T}^{\equiv 1}_{3,2}$. So, $w(\texttt{T}^{=}_{2}\setminus\texttt{U}^{=}_{2})\geq 2p_{ch}^{0}+2u_{co}$.
@@
    $w(\texttt{T}^{\neq}_{2})=\sum_{i\in[n],j\in[11k]:j\nmid 11}\kappa_{i,j}.$
    Since $T$ was a valid tour, we have $\kappa_{i,j}\geq 2$ for all such $i,j$. Furthermore, if either of the involved vertices is dishonest, the fact that the degree of that vertex in $T$ is even implies that $\kappa_{i,j}$ is odd. As a consequence, we have $\kappa_{i,j}\geq 3$ for at least $d_{ch}/2$ such $(i,j)$. This implies $w(\texttt{T}^{\neq}_{2})\geq 2\cdot 10nk+d_{ch}/2$.
  • 3LIN(2) clauses: Let $\zeta=\texttt{P}^{\equiv 1}_{3,2}(x_{i_{1},j_{1},b_{1}},x_{i_{2},j_{2},b_{2}},x_{i_{3},j_{3},b_{3}})$ be a clause in $\texttt{I}^{\equiv 1}_{3,2}$. Let $T_{\zeta}$ be the set of edges in $\texttt{T}^{\equiv 1}_{3,2}$ involved in clause $\zeta$. We will extract a valid spanning tour $Q$ of $H$ from $T_{\zeta}$, and then use the soundness property of $H$ to argue that $w(T_{\zeta})$ is large.
@@
    Summing over all $2nk/3$ clauses $c$ in $\texttt{I}^{\equiv 1}_{3,2}$, we get
    $w(\texttt{T}^{\equiv 1}_{3,2})\geq(2nk/3)\cdot s(H)-2u_{co}-t_{co}+r,$
@@
Together, these imply the lower bound $w(T)=w(\texttt{U}^{=}_{2})+w(\texttt{T}^{=}_{2})+w(\texttt{T}^{\neq}_{2})+w(\texttt{T}^{\equiv 1}_{3,2})\geq nk\cdot(30+2\cdot s(H)/3)+(d_{ch}+d_{co})/2+p_{ch}+r$. ∎
@@ D.2 Description of our gadget
We note that this is in alignment with the discussion in Section 4; every satisfying assignment to $\texttt{P}^{\equiv 1}_{3,2}$ admits a tour of cost at most $c(H)=10$, whereas for every tour $Q$ corresponding to an unsatisfying assignment, we have $k_{1}^{H}(Q)=0$ and $k_{2}^{H}(Q)$ is even; hence $Q$ has weight at least $s(H)+\mathbf{1}\{k_{1}^{H}(Q)=0\}\cdot\mathbf{1}\{k_{2}^{H}(Q)\text{ is even}\}=11$. We can now prove Theorem 4.1.